# Rogue waves in extended Gross-Pitaevskii Models with a Lee-Huang-Yang correction

### v2

@@ II Extended model &\& its reduction
The system under investigation consists of a 1D homonuclear bosonic mixture. Here, atoms of, e.g., 39K reside in two different hyperfine states, according to the corresponding 3D setup of Semeghini et al. (2018) in free space. In the 1D case, the corresponding geometry is achieved by utilizing tightly confined transversal directions as compared to the elongated unconfined longitudinal direction. Both states are assumed to share the same atom number ($N_{1}=N_{2}\equiv N$), feature equal intra-component repulsion, $g_{11}=g_{22}\equiv g>0$, and inter-component attraction, $g_{12}<0$, yielding a 1D droplet-bearing environment for $\delta g=g_{12}+g>0$. Under these assumptions, the two components become equivalent, allowing the mixture to be described by a reduced single-component 1D eGPE Petrov (2015); Petrov and Astrakharchik (2016), incorporating the leading-order LHY quantum correction. In dimensionless form, this model reads:
@@
where $\psi(x,t)$ denotes the 1D wavefunction, $\rho(x,t)=|\psi|^{2}$ is the local density, and $f(\rho)=-\rho+\sqrt{\rho}$ contains the nonlinear interaction terms. The quadratic nonlinearity captures the effectively attractive, 1D contribution of the LHY correction, while the cubic term represents the standard mean-field repulsion. Finally, in Eq. (1), energy is measured in units of $\hbar^{2}/(m\xi^{2})$, with $\xi=\pi\hbar^{2}\sqrt{|\delta g|}/(mg\sqrt{2g})$ denoting the healing length and $m$ is the atom mass. Also, time, length, and wavefunction are measured in units of $\hbar/\left(m\xi^{2}\right)$, $\xi$ and $(2\sqrt{g})^{3/2}/(\pi\xi(2|\delta g|)^{3/4})$ respectively.
@@ III Existence of spatiotemporal RWs
Unlike the integrable cubic NLS/GPE, the eGPE does not —to the best of our knowledge— admit explicit analytical formulas for rational solutions or space-time periodic RW-like generalizations thereof. Therefore, alternative numerical computation-based methods are necessary to establish their existence.
@@
where $\tilde{t}=t-\frac{T}{2}$ and $\alpha_{0}=\rho_{0}f^{\prime}(\rho_{0})=-\rho_{0}+\sqrt{\rho_{0}}/2$ is the dilatation factor. Here, we emphasize that although Eq. (4) is not an asymptotically valid approximation of the eGPE RW, it nevertheless remains sufficiently accurate for small $|\mu_{0}|\ll 1$ serving as a useful approximation or as an initial guess for seeding the fixed-point computations of the boundary value problem.
@@
Figure 1: Representative eGPE RW moduli profiles, $|\psi(x,t)|$. (a) Space-time profile of a rogue waveform for $\mu_{0}\approx-0.12$, and (b) a second-order rogue waveform characterized by $\mu_{0}\approx-0.02$. (c) [(d)] Snapshot of the eGPE [second-order] rogue waveform in (a) [(b)] at $\tilde{t}=0$ compared to a cubic GPE PS [second-order RW] possessing the same peak amplitude. Parameters used for the PS [HORW] are $L=30$ and $T=190$ [$L=60$ and $T=290$].
@@
Figure 2: RW families of the eGPE. The $\phi||_{2}$ norm of the spatiotemporally identified fixed point $\phi$ starting from (a) an NLS-like PS (first family) and (h) a sech-like waveform (second family) as a function of the chemical potential $\mu_{0}$. Inset in (h) depicts a magnification of $\phi||_{2}$ at $\mu_{0}\approx-0.1935$, see the green rectangle. Representative waveforms $|\phi(x,t)|$ of the upper (lower) solution branch (i.e., curve) depicted in panel (a) with pink (purple) are shown in panels (b) -(d) ((e)-(g)). Here, panel (b) [(d)] depicts a configuration with $\mu_{0}\approx-0.1009$ $\mu_{0}\approx-0.2$] and panel (e) [(g)] refers to $\mu_{0}\approx-0.2$ $\mu_{0}\approx-0.09$]. We remark that the purple line (hardly visible) near the turning point at $\mu_{0}\approx-0.2$ practically coincides with the pink. Note the progressive deformation from PS-like structures belonging to the upper branch towards wider RW patterns as $\mu_{0}$ decreases, and from multi-hump structures (e.g., in (f)) towards PS-like entities (in (g)) in the lower branch as $\mu_{0}$ increases. In the second RW family, panel (h), we observe a bifurcation from a homogeneous background in (i) to an NLS periodic wave in (j) and gradually a PS-like solution for decreasing $\mu_{0}$ in (k). Multi-hump RWs occur for further decreasing values of $\mu_{0}$ (panels (l)-(n)). Vertical dashed lines indicate the values of $\mu_{0}$ for the solutions depicted in panels (b)-(g) and (i)-(n).
To trace families of RW-like solutions we start from a numerical eGPE RW (converged, upon iterative usage of the NLS approximation) and subsequently perform parametric continuation over the chemical potential $\mu_{0}$. Our findings are summarized in Fig. 2. Specifically, panels (a) and (h) within Fig. 2 present the branches of the two distinct families of RW solutions identified herein. An appropriate matrix norm, $\phi||_{2}=\bar{\sigma}=\underset{i=1,2,\cdots}{\rm max}[\sigma_{i}]$,where ${\sigma}_{i}$ represents the singular value decomposition of the discretized $\phi$ (the $L^{2}$ norm –see, e.g., Ref. Quarteroni et al. (2010)), of each computed solution, is presented against $\mu_{0}$, yielding a relevant bifurcation diagram.
@@
For intermediate values of $\mu_{0}\approx-0.1005$, the central hump splits along the time axis. As such, this doubly periodic waveform coexists with a large amplitude periodic wave background –see Fig. 2(l). With decreasing (more negative) chemical potential, we approach the first turning point ($\mu_{0}\approx-0.1934$) of the existence curve. The solutions corresponding to these chemical potentials exhibit significant mass concentration in their central humps (Fig.2(n)). Extending the continuation beyond the first turning point reveals a second turning point in the bifurcation structure occuring at $\mu_{0}\approx-0.1933$ (see inset of Fig. 2(h)). We remark that the qualitative features of the relevant solutions remain unchanged across this second fold (results not shown here for brevity). Having identified the existence of two distinct families of eGPE RWs, namely the first (second) family shown in Fig 2 (b)-(g) (Fig 2 (i)-(n)), we next aim to inspect their dynamical stability by direct spatio-temporal evolution. Figure 3(a) presents the evolution of a representative single-hump member (Fig. 2(c)) of the first family of solutions. As can be seen, it substantially deforms after two temporal periods. This behavior is a general feature of single hump solutions spanning both families. In contrast, robust propagation is observed (see Fig. 3(b)) for the two-hump structure depicted in Fig. 2(l). Evidently, such states survive for several periods before shape deformations manifested in their tails and core take place. Similar dynamics generically pertains to multihump solutions within both families. More elaborate studies on the (in)stability of these (short-) long-lived (single-) multi-hump solutions in terms of Floquet theory DECONINCK and OLIVERAS (2011); Cuevas-Maraver et al. (2017) is deferred to a future work.
@@ IV.1 Interference of dam breaks
as an initial guess together with the computational parameters $L\approx 92.69$ and $T=200$. Here, the AB breather is characterized by two free parameters. Namely, the density $\rho_{0}$ and the control parameter [^4] Tuning $a$ towards $0.5$ leads to the PS. $0<a<0.5$, while $\omega=2\sqrt{1-2a}$ and $b=\sqrt{8a(1-2a)}$ are determined by $a$. Note also the excellent agreement between the central lattice element (i.e., centered about $x=0$) and a numerically identified PS-like wave obtained from the boundary value problem with $\mu_{0}\approx-0.0767$, $L=40$, and $T=190$ (Fig. 4(c)). This observation signals the progressive deformation and in particular, the lengthening of the spatial period of the modulated RW lattice El et al. (2016).
Additionally, the dynamical formation of the RW lattice configuration is accompanied by the emission of counterpropagating matter wave jets. These are a manifestation of the self-evaporation Ferioli et al. (2020) of the droplet background due to its excited nature induced by the interference process. Such self-evaporation mechanism, leading to density emission from a 1D droplet core has also been argued to occur after droplets collision Astrakharchik and Malomed (2018); Katsimiga et al. (2023a) and in the presence of a dynamically switched-on defect Bristy et al. (2025). It is worth mentioning at this point that an overall similar response takes place within the standard GPE model (not shown) but without the above-discussed emission of droplet-like density portions. This further supports the importance of the LHY contribution and marks the differences of the phenomenology in the presence of competing nonlinear interactions.
@@
Figure 5: (a) Interfering DBFs for $\rho_{0}=0.0875$ leading to RW generation. (b) Snapshot of $|\psi|$ features an eGPE RW compared with the computed rogue waveform shown in Fig. 2(a) having $\mu_{0}\approx-0.1713$. (c) A structure reminiscent of a HORW. (d) Snapshot illustrates a bimodal breathing state, which subsequently merges into (e) a larger amplitude structure centered at $x=0$. Vertical dashed lines in panel (a) mark the time-instants presented in panels (b)-(e).
@@ IV.2 Gradient Catastrophe
The underlying gradient catastrophe mechanism within this attractive regime leads to the formation of a rogue event, at $t=84$ (Fig. 6 (c)). This is compared to a numerically computed eGPE RW, shown as a yellow dashed line, obtained via a fixed point computation with $\mu_{0}=-0.1138$ (see Fig. 2(a)), demonstrating very good agreement. At $t=160$, an apparent HORW event emerges, which then recurs —with the nodal points expanding outwards— at $t=320$ (Figs. 6(d),(e)). Over longer times, a highly coherent nearly time-periodic excitation forms as can be seen by the wavefunction amplitude snapshot presented in Fig. 6(f); this is again reminiscent of two-soliton breathing solutions within the integrable, cubic NLS Satsuma and Yajima (1974).
@@ Acknowledgements.
The authors are thankful to M. A. Hoefer and E. G. Charalampidis for insightful discussions. This material is based upon work supported by the US National Science Foundation under Grant No. PHY- 2110030, PHY-2408988 and DMS-2204702 (P.G.K.). S.I.M. acknowledges support from the Missouri University of Science and Technology, Department of Physics, Startup fund.

### v3

@@ II Extended model &\& its reduction
The system under investigation consists of a 1D homonuclear bosonic mixture. Here, atoms of, e.g., 39K reside in two different hyperfine states, according to the corresponding 3D setup of Semeghini et al. (2018) in free space. In the 1D case, the corresponding geometry is achieved by utilizing tightly confined transversal directions as compared to the elongated unconfined longitudinal direction. Both states are assumed to share the same atom number ($N_{1}=N_{2}\equiv N$), feature equal intra-component repulsion, $g_{11}=g_{22}\equiv g>0$, and inter-component attraction, $g_{12}<0$, yielding a 1D droplet-bearing environment for $\delta g=g_{12}+g>0$. Under these assumptions, the two components become equivalent, allowing the mixture to be described by a reduced single-component 1D eGPE Petrov (2015); Petrov and Astrakharchik (2016), incorporating the leading-order LHY quantum correction. Notably, the latter is attractive in 1D Petrov and Astrakharchik (2016) in sharp contrast to its repulsive 3D counterpart Petrov (2015). In dimensionless form, this model reads:
@@
where $\psi(x,t)$ denotes the 1D wavefunction, $\rho(x,t)=|\psi|^{2}$ is the local density, and $f(\rho)=-\rho+\sqrt{\rho}$ contains the nonlinear interaction terms. The quadratic nonlinearity captures the effectively attractive, 1D contribution of the LHY correction, while the cubic term represents the standard mean-field repulsion. Finally, in Eq. (1), energy is measured in units of $\hbar^{2}/(m\xi^{2})$, with $\xi=\pi\hbar^{2}\sqrt{|\delta g|}/(mg\sqrt{2g})$ denoting the healing length and $m$ is the atom mass. Also, time, length, and wavefunction are measured in units of $\hbar/\left(m\xi^{2}\right)$, $\xi$ and $(2\sqrt{g})^{3/2}/(\pi\xi(2|\delta g|)^{3/4})$ respectively. Representative dimensionless evolution times considered herein correspond to $t\sim 10^{3}$, which refer to $\sim 245$ms in dimensional units, e.g., for a transverse trap frequency of $\sim 2\pi\times 650$Hz proximal to the droplet experiment of Ref. Cabrera et al. (2018). Recall that we assume no longitudinal trap herein.
@@ III Existence of spatiotemporal RWs
Unlike the integrable cubic NLS (or GPE), the eGPE does not —to the best of our knowledge— admit explicit analytical formulas for rational solutions or space-time periodic RW-like generalizations thereof. Therefore, alternative numerical computation-based methods are necessary to establish their existence.
@@
where $\tilde{t}=t-\frac{T}{2}$ and $\alpha_{0}=\rho_{0}f^{\prime}(\rho_{0})=-\rho_{0}+\sqrt{\rho_{0}}/2$ is the dilatation factor. Here, we emphasize that although Eq. (4) is not an asymptotically valid approximation of the eGPE RW, it nevertheless remains sufficiently accurate for small $|\mu_{0}|\ll 1$ serving as a useful approximation or as an initial guess for seeding the fixed-point computations of the boundary value problem. Thus, in the small $|\mu_{0}|$ limit, the eGPE RW is anticipated to resemble its integrable counterpart known from the cubic NLS El et al. (2016); Bertola and Tovbis (2013), as we argue below. This fact partially highlights the robustness of NLS RWs in non-integrable systems and reflects the generic nature of the relevant mechanisms (going beyond the specific case example of integrable systems).
@@
Figure 1: Representative eGPE RW moduli profiles, $|\psi(x,t)|$. (a) Space-time profile of a rogue waveform for $\mu_{0}\approx-0.12$, and (b) a second-order rogue waveform characterized by $\mu_{0}\approx-0.02$. (c) [(d)] Snapshot of the eGPE first-order [second-order] rogue waveform in (a) [(b)] at $\tilde{t}=0$ compared to a cubic GPE PS [second-order RW] possessing the same peak amplitude. Parameters used for the PS [HORW] are $L=30$ and $T=190$ [$L=60$ and $T=290$].
@@
Figure 2: RW families of the eGPE. The $\phi||_{2}$ norm of the spatiotemporally identified fixed point $\phi$ starting from (a) an NLS-like PS (first family) and (h) a sech-like waveform (second family) as a function of the chemical potential $\mu_{0}$. Inset in (h) shows a magnification of $\phi||_{2}$ at $\mu_{0}\approx-0.1935$, see the green rectangle. Representative waveforms $|\phi(x,t)|$ of the upper (lower) solution branch (i.e., curve) depicted in panel (a) with pink (purple) are shown in panels (b) -(d) ((e)-(g)). Here, panel (b) [(d)] depicts a configuration with $\mu_{0}\approx-0.1009$ $\mu_{0}\approx-0.2$] and panel (e) [(g)] refers to $\mu_{0}\approx-0.2$ $\mu_{0}\approx-0.09$]. We remark that the purple line (hardly visible) near the turning point at $\mu_{0}\approx-0.2$ practically coincides with the pink. Note the progressive deformation from PS-like structures belonging to the upper branch towards wider RW patterns as $\mu_{0}$ decreases, and from multi-hump structures (e.g., in (f)) towards PS-like entities (in (g)) in the lower branch as $\mu_{0}$ increases. In the second RW family, panel (h), we observe a bifurcation from a homogeneous background in (i) to an NLS periodic wave in (j) and gradually a PS-like solution for decreasing $\mu_{0}$ in (k). Multi-hump RWs occur for further decreasing values of $\mu_{0}$ (panels (l)-(n)). Vertical dashed lines indicate the values of $\mu_{0}$ for the solutions depicted in panels (b)-(g) and (i)-(n).
To trace families of RW-like solutions we start from a numerical eGPE RW (converged, upon iterative usage of the NLS RW approximation) and subsequently perform parametric continuation over the chemical potential $\mu_{0}$. Our findings are summarized in Fig. 2. Specifically, panels (a) and (h) within Fig. 2 present the branches of the two distinct families of RW solutions identified herein. An appropriate matrix norm, $\phi||_{2}=\bar{\sigma}=\underset{i=1,2,\cdots}{\rm max}[\sigma_{i}]$,where ${\sigma}_{i}$ represents the singular value decomposition of the discretized $\phi$ (the $L^{2}$ norm –see, e.g., Ref. Quarteroni et al. (2010)), of each computed solution, is presented against $\mu_{0}$, yielding a relevant bifurcation diagram.
@@
For intermediate values of $\mu_{0}\approx-0.1005$, the central hump splits along the time axis. As such, this doubly periodic waveform coexists with a large amplitude periodic wave background –see Fig. 2(l). With decreasing (more negative) chemical potential, we approach the first turning point ($\mu_{0}\approx-0.1934$) of the existence curve. The solutions corresponding to these chemical potentials exhibit significant mass concentration in their central humps (Fig.2(n)). Extending the continuation beyond the first turning point reveals a second turning point in the bifurcation structure occuring at $\mu_{0}\approx-0.1933$ (see inset of Fig. 2(h)). We remark that the qualitative features of the relevant solutions remain unchanged across this second fold (results not shown here for brevity).
Having identified the existence of two distinct families of eGPE RWs, namely the first [second] family shown in Fig 2 (b)-(g) (Fig 2 (i)-(n)), we next aim to inspect their dynamical stability by direct spatio-temporal evolution. Figure 3(a) presents the evolution of a representative single-hump member (Fig. 2(c)) of the first family of solutions. As can be seen, it substantially deforms after two temporal periods. This behavior is a general feature of single hump solutions spanning both families. In contrast, robust propagation is observed (see Fig. 3(b)) for the two-hump structure depicted in Fig. 2(l). Evidently, such states survive for several periods before shape deformations manifested in their tails and core take place. Similar dynamics generically pertains to multihump solutions within both families. More elaborate studies on the (in)stability of these (short-) long-lived (single-) multi-hump solutions in terms of Floquet theory DECONINCK and OLIVERAS (2011); Cuevas-Maraver et al. (2017) is deferred to a future work.
@@ IV.1 Interference of dam breaks
as an initial guess together with the computational parameters $L\approx 92.69$ and $T=200$. Here, the AB breather is characterized by two free parameters. Namely, the background density $\rho_{0}$ and the control parameter [^4] Tuning $a$ towards $0.5$ leads to the PS. $0<a<0.5$, while $\omega=2\sqrt{1-2a}$ and $b=\sqrt{8a(1-2a)}$ are determined by $a$. Note also the excellent agreement between the central lattice element (i.e., centered about $x=0$) and a numerically identified PS-like wave obtained from the boundary value problem with $\mu_{0}\approx-0.0767$, $L=40$, and $T=190$ (Fig. 4(c)). This observation signals the progressive deformation and in particular, the lengthening of the spatial period of the modulated RW lattice El et al. (2016).
Additionally, the dynamical formation of the RW lattice configuration is accompanied by the emission of counterpropagating matter wave jets. These are a manifestation of the self-evaporation Ferioli et al. (2020) of the droplet background due to its excited nature induced by the interference process. Such self-evaporation mechanism, leading to density emission from a 1D droplet core has also been argued to occur after droplet collision Astrakharchik and Malomed (2018); Katsimiga et al. (2023a) and in the presence of a dynamically switched-on defect Bristy et al. (2025). It is worth mentioning at this point that an overall similar response takes place within the standard GPE model (not shown) but without the above-discussed emission of droplet-like density portions. This further supports the importance of the LHY contribution and marks the differences of the phenomenology in the presence of competing nonlinear interactions.
@@
Figure 5: (a) Interfering DBFs for $\rho_{0}=0.0875$ leading to RW generation. (b) Snapshot of $|\psi|$ featuring an eGPE RW which is compared with the computed rogue waveform shown in Fig. 2(a) having $\mu_{0}\approx-0.1713$. (c) A structure reminiscent of a HORW. (d) The snapshot illustrates a bimodal breathing state, which subsequently merges into (e) a larger amplitude structure centered at $x=0$. Vertical dashed lines in panel (a) mark the time-instants presented in panels (b)-(e).
@@ IV.2 Gradient Catastrophe
The underlying gradient catastrophe mechanism within this attractive regime leads to the formation of a rogue event, at $t=84$ (Fig. 6 (c)). This is compared to a numerically computed eGPE RW, shown as a yellow dashed line, obtained via a fixed point computation with $\mu_{0}=-0.1138$ (see Fig. 2(a)), demonstrating very good agreement. At $t=160$, an event reminiscent of a HORW emerges, which then recurs —with the nodal points expanding outwards— at $t=320$ (Figs. 6(d),(e)). Over longer times, a highly coherent nearly time-periodic excitation forms as can be seen by the wavefunction amplitude snapshot presented in Fig. 6(f); this is again reminiscent of two-soliton breathing solutions within the integrable, cubic NLS Satsuma and Yajima (1974).
@@ Appendix A Dam breaks in the attractive regime
Appendix B DATA AVAILABILITY
The data that support the findings of this article are not publicly available upon publication because it is not technically feasible and/or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.
@@ Acknowledgements.
The authors are thankful to M. A. Hoefer and E. G. Charalampidis for insightful discussions. This material is based upon work supported by the US National Science Foundation under Grant No. PHY- 2110030, PHY-2408988 and DMS-2204702 (P.G.K.). S.I.M. acknowledges support from the Missouri University of Science and Technology, Department of Physics, Startup fund. This research was partly conducted while P.G.K. was visiting the Okinawa Institute of Science and Technology (OIST) through the Theoretical Sciences Visiting Program (TSVP). This work was also supported by the Simons Foundation [SFI-MPS-SFM-00011048, P.G.K.].